Calculate the molarity of the following solutions.

A.) 0.800 mole glucose in 0.450 L of a glucose solution

Express your answer with the appropriate units.

B.) 83.5 g of HCl in 2.00 L of a HCl solution

Express your answer with the appropriate units.

C.) 36.5 g of NaOH in 350. mL of a NaOH solution

Express your answer with the appropriate units.
Answer #1

A)

volume , V = 0.45 L

use:

Molarity,

M = number of mol / volume in L

= 0.8/0.45

= 1.778 M

Answer: 1.18 M

B)

Molar mass of HCl,

MM = 1*MM(H) + 1*MM(Cl)

= 1*1.008 + 1*35.45

= 36.458 g/mol

mass(HCl)= 83.5 g

use:

number of mol of HCl,

n = mass of HCl/molar mass of HCl

=(83.5 g)/(36.46 g/mol)

= 2.29 mol

volume , V = 2 L

use:

Molarity,

M = number of mol / volume in L

= 2.29/2

= 1.145 M

Answer: 1.15 M

C)

Molar mass of NaOH,

MM = 1*MM(Na) + 1*MM(O) + 1*MM(H)

= 1*22.99 + 1*16.0 + 1*1.008

= 39.998 g/mol

mass(NaOH)= 36.5 g

use:

number of mol of NaOH,

n = mass of NaOH/molar mass of NaOH

=(36.5 g)/(40 g/mol)

= 0.9125 mol

volume , V = 3.5*10^2 mL

= 0.35 L

use:

Molarity,

M = number of mol / volume in L

= 0.9125/0.35

= 2.607 M

Answer: 2.61 M